OpenStudy (anonymous):

how do i completely factor this?? x^2-16x+55

pooja195 (pooja195):

Forget about the = 0 for now. (x-11)(x-5) is factored. If you want to solve: (x-11)(x-5)=0 Set each equal to 0. x-11=0 x-5=0 x=11 x=5 So x=5,11

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

np

OpenStudy (anonymous):

would you be abel to help me on my other math homework and help me show my work bc if i dont my teacher will yell at me

pooja195 (pooja195):

suree

OpenStudy (anonymous):

now i got to factor trinomials 6x^2+17x+12

pooja195 (pooja195):

There are a few ways to do this, but the easiest way I know of to factor a trinomial in the form ax^2 + bx + c is the following: 1. Multiply a and c. 2. If the sign of c is positive, select a pair of factors for the product ac that will have a sum of b. If the sign of c is negative, select a pair of factors for ac that will have a difference of b. 3. Replace the term bx with either the sum or difference. 4. Group and factor. In this problem, 6*12 = 72. Since it is positive 12, we need a pair of factors of 72 that has a sum of -17. This pair of factors is -8 and -9. So we rewrite the original expression as 6x^2 - 9x - 8x + 12. Then, separate this expression into two groups of two terms using parentheses. (6x^2 - 9x) + (-8x + 12) Within each pair of parentheses, factor out any common factor. 3x(2x - 3) + -4(2x - 3) Notice that both parenthetical expressions are the same. Factor out this common parenthetical expression. The result is (2x - 3) (3x - 4)

OpenStudy (anonymous):

doing the same 6y^2+7y-24

pooja195 (pooja195):

6y² + 7y - 24 = 0 y² + 7/12y = 4 + (7/12)² y² + 7/12y = 576/144 + 49/144 (y + 7/12)² = 625/144 y + 7/12 = +/- 25/12 Factors: = y + 7/12 - 25/12, = y - 18/12, = y - 3/2, = 2y - 3 = y + 7/12 + 25/12, = y + 32/12, = y + 8/3, = 3y + 8 Answer: (2y - 3)(3y + 8) Checking: = (2y - 3)(3y + 8) = 6y² + 16y - 9y - 24 = 6y² + 7y - 2

OpenStudy (anonymous):

can we do three more plzz bcnow i think i know how to do it

OpenStudy (anonymous):

16r^2-16r-12

OpenStudy (anonymous):

4x^2-20x+25

OpenStudy (anonymous):

4x^2+7x+3

pooja195 (pooja195):

hold on

pooja195 (pooja195):

@jim_thompson5910

OpenStudy (anonymous):

whats that for

jimthompson5910 (jim_thompson5910):

16r^2-16r-12 4(4r^2-4r-3) ... factor out the GCF 4 4(4r^2+2r-6r-3) ... break up -4r into 2r - 6r (see note below) 4((4r^2+2r)+(-6r-3)) 4(2r(2r+1)+(-6r-3)) 4(2r(2r+1)-3(2r+1)) 4(2r-3)(2r+1) --------------------------------------------------------------------------- 16r^2-16r-12 completely factors to 4(2r-3)(2r+1)

jimthompson5910 (jim_thompson5910):

Note: the numbers 2 and -6 both add to -4 AND multiply to -12

jimthompson5910 (jim_thompson5910):

so that explains why I rewrote 4r^2-4r-3 into 4r^2+2r-6r-3

jimthompson5910 (jim_thompson5910):

let me know if anything doesn't make any sense

OpenStudy (anonymous):

it is

jimthompson5910 (jim_thompson5910):

so it all makes sense?